OpenStudy (anonymous):

For an open-topped rectangular box, with a square base x by x cm and height h cm, find the dimensions giving the minimum surface area, given that the volume is 13 cm^3. ***write EXACT answers*** x=_____cm h=_____cm please explain! Thank you!! :)

OpenStudy (johnweldon1993):

Well this is definitely...odd, lets see what we can do here...

OpenStudy (johnweldon1993):

Well for a rectangular prism...the volume is going to be LxWxH L and W are the same 'x' ...so lets write \[\large V = x^2 \times h\]

OpenStudy (anonymous):

yeah haha definitely odd :P okay i'm following so far:)

OpenStudy (anonymous):

so then you go \[13=x^2 * h\] ?

OpenStudy (johnweldon1993):

Well we're going to want the area of the bottom...plus the 4 sides of the box (4 sides because we have the base...and then the 4 sides (open top))

OpenStudy (anonymous):

okay... so multiply by 4?

OpenStudy (johnweldon1993):

So first lets express that Volume equation in terms of h if V = x^2 x H H = V/x^2

OpenStudy (anonymous):

okay:)

OpenStudy (johnweldon1993):

Alright...so the area of the bottom...is just the area of a square..so Length x Width....which we know is x^2 right?

OpenStudy (anonymous):

mhmm:)

OpenStudy (johnweldon1993):

But then we need to ADD the area of the 4 sides... Well each of the 4 sides have an area of (x times h) So...altogether we have A = x^2 + (4xh) We know h = v/x^2 so \[\large A = x^2 + 4x(\frac{V}{x^2})\]

OpenStudy (johnweldon1993):

Alright yeah I think I'm coming out with the right answer lol..okay have faith we're almost done haha

OpenStudy (anonymous):

hahaha of course :P okay, so i followed that part:)

OpenStudy (johnweldon1993):

Alright.. well we can see that an 'x' can be canceled there on the right \[\large A = x^2 + \frac{4V}{x}\] follow that?

OpenStudy (anonymous):

yes :)

OpenStudy (johnweldon1993):

Awesome... okay here's the only annoying part "minimize the surface area" -_- haha not too bad! We just need to take the derivative of this function and set it equal to 0

OpenStudy (anonymous):

okay yeah this stuff is annoying haha ermm so A' = 2x - 4V ---- x^2 ? not sure about that second term :/

OpenStudy (johnweldon1993):

How are you not sure!? That's completely correct! :) lol

OpenStudy (anonymous):

oh!! okay awesome!! yay :) hehe so \[2x - \frac{ 4V }{ x^2 } = 0\] ?

OpenStudy (anonymous):

are we solving for x?

OpenStudy (johnweldon1993):

Alright...so remember that is equal to 0 \[\large 2x - \frac{4V}{x^2} = 0\] So we can do \[\large 2x = \frac{4V}{x^2}\] and yes solving for 'x' there

OpenStudy (anonymous):

okay, so 2x^3 = 4V x^3 = 2V ? so x = cube root 2V ? ;/

OpenStudy (johnweldon1993):

Perfect again...haha you and your "unsure face" :P

OpenStudy (anonymous):

haha woo! and yeah.... it's appropriate though since i wasn't sure ahhahaa :P

OpenStudy (johnweldon1993):

So looky we have our 2 equations to solve \[\huge x = \sqrt[3]{2V}\] and \[\huge h = \frac{V}{x^2}\] so plug in your 13 for V and solve for 'x' and then you can solve for 'h'

OpenStudy (anonymous):

okay:) ermm so x= cube root 2(13) =cube root 26 ?? would that be considered the exact answer for x? :/

OpenStudy (johnweldon1993):

Indeed! they want to be PICKY and be EXACT! haha

OpenStudy (anonymous):

ahh okay! :) so x= cube root 26 and h=13/(cube root 26)^2 ?? and would we just leave that like that?

OpenStudy (johnweldon1993):

Yeah...I mean that is correct \[\huge x = \sqrt[3]{26}\] \[\huge h = \frac{13}{\sqrt[3]{26}^2}\] I 'm not sure if they want 'h' like that or like \[\huge h = \frac{13}{26^{\frac{2}{3}}}\] but yeah...either way ...you arrive at the correct solution

OpenStudy (anonymous):

ahh okay:) yay!! :) so that's all there is to this problem? :O

OpenStudy (johnweldon1993):

That's it.. and we can check obviously by doing \[\huge V = x^2 \times h\] \[\huge 13 = \sqrt[3]{26}^2 \times \frac{13}{\sqrt[3]{26}^2}\] \[\huge 13 = \frac{13\times \cancel{26^{2/3}}}{\cancel{26^{2/3}}} = 13\] hey looky there :D

OpenStudy (anonymous):

haha yeah looky there! :) awesome!! Thanks so much!! :D this makes sense now!! :D

OpenStudy (johnweldon1993):

No problem :)